Radon
-
Important operating system updates - Community Clusters
On the morning of Thursday, February 5, 2015, Carter, Conte, Hansen, Peregrine1, Radon, and Rossmann login servers will be rebooted to apply an important Red Hat Linux operating system update. Additionally, during this time scratch storage servers w...
-
Security updates for Conte and Radon scratch
Conte scratch has returned to full production as of approximately 2:15 pm. Update: February 26, 2015 1:20pm Radon scratch has returned to full production as of approximately 10:30 am. The work on Conte scratch is progressing as planned. Original mes...
-
Research Data Depot Security Updates
As of 3:15 pm the maintenance is complete and Research Data Depot is returned to full production. Original message: The storage servers powering the Research Data Depot will undergo maintenance on Thursday, February 26, 2015 from 10:00am - 4:00pm EST...
-
The Radon cluster will be unavailable on Tuesday, 17 March, 2015. On that day, Radon's job scheduler will be updated, and its network connection to its scratch system reconfigured. As of 5:30pm Tuesday, March 17, Radon cluster has been returned to se...
-
The Hammer, Hathi, Radon, and Snyder cluster will be unavailable beginning at Wednesday, July 1, 2015 from 8:00am - 12:00pm EDT, for scheduled maintenance. The cluster will return to full production by Wednesday, July 1st, 2015 at 12:00pm EDT. The do...
-
Due to power work in the MSEE building, most ECN services will be unavailable between 6:30am – 9:00pm EDT on Saturday, August 15, 2015. For Research Computing users this means that software packages licensed through ECN servers will not be able to ch...
-
November 3, 2015 6:15pm The maintenance for Radon is completed and the cluster has been returned to production. Original The Radon cluster will be unavailable beginning at Tuesday, November 3, 2015 from 7:00am - 7:00pm EST, for scheduled maintenance....
-
Unscheduled Home Filesystem Outage
As of 12:46, December 2, the home filesystem serving Conte, Hammer, Hansen, Hathi, Peregrine1, Radon, Rice, and Snyder was restored to normal operation. All queues have been re-enabled. As of Wednesday, December 2nd, 2015 at 12:00pm EST, Conte, Hamm...
-
Unscheduled Outage in Math Data Center
Most of the impact of this turned out to be to the Depot storage system, which has now been restored to normal operations. All the other affected systems are showing a return to normal operations now. Original Message: As of Thursday, February 4th,...
-
ECN services outage - ITaP Research Computing systems impacted
Engineering Computing Network (ECN) will be performing staged patching and reboots of all of ECN's RedHat Linux workstations and servers to protect against a serious vulnerability in glibc system library. A significant number of ECN services will be...
-
The Isilon filesystem was restored to normal service and all affected clusters had it remounted as quickly as was sustainable by the filesystem. This process was completed by Wednesday, May 18th, 2016 at 12:15am EDT. All clusters other than Conte (...
-
An issue was discovered shortly after Conte, Hansen, Hathi, and Radon were brought back online with the /group path on several front-ends and nodes. Any scripts or jobs that rely on the /group path may have had issues immediately following return to...
-
The underlying storage has been fixed, and all these clusters have been returned to normal operations as of 10:00pm EDT. As of Tuesday, June 7th, 2016 at 4:10pm EDT, Conte, Hansen, Hathi, and Radon are unavailable due to a loss of Isilon home direct...
-
Engineering Computing Network (ECN) will be performing scheduled maintenance this weekend on several ECN server resulting in their unavailability for a short time. Some ECN services will be affected, including several software license servers for ITa...
-
Update As of 5:50 pm, Tuesday, 16 Aug 2016, the Radon cluster has been returned to service and is fully operational. Thank you for your patience. Update Due to unanticipated conflicts between the upgraded scheduler and our network configuration, the...
-
Degraded performance of several systems
We have seen a significant wave of these events this morning, September 21. For the most part, this wave seems to have been linked to a storage problem that has been resolved. However, we are implementing new monitoring and response procedures toda...
-
Home Filesystem Maintenance - All Clusters
Conte has been returned to normal operations as well now. This concludes the home directory maintenance on all systems. Update: September 27, 2016 11:55pm All systems other than Conte have been successfully returned to normal operations with the ne...
-
Job scheduling paused on Radon
Job scheduling was paused on Radon between 6 pm and 7 pm this evening. Node monitoring processes marked most nodes offline around 6 pm, preventing new jobs from starting. System engineers cleared the fault in the node monitoring, and nodes came back...
-
Emergency Security Patching of RCAC Clusters
Due to a recent security vulnerability, the Carter, Halstead, Hammer, Radon, Rice, Scholar, and Snyder clusters will have their operating system upgraded to a newer version during February 2, 2017 5:00pm - March 2, 2017 5:00pm EST. Unlike other cl...
-
Engineers have restored failed core servers back to a functional state. Data Depot is up and running as normal and job scheduling resumed. Should you encounter any lingering issues please let us know at rcac-help@purdue.edu Original Message Some core...